The most recent financial statements for Cardinal, Inc., are shown here: Income Statement Sales Costs Balance Sheet $29,800 Assets $70,800 Debt $34,600 Equity 36,200 18,250 $11,550 Total $70,800 Total $70,800 income Taxes (22%) 2541 Net income $ 9,009 Assets and costs are proportional to sales. Debt and equity are not. A dividend of $3,400 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to be $33,972. What is the external financing needed? (Do not round intermediate calculations.) External financing needed
Answer #1

Growth rate in sales=(33972-29800)/29800=14%

Dividend payout ratio=Dividend/Net income

=(3400/9009)=0.377400377

Sales 33972
Costs(18250*1.14) 20805
Taxable income $13167
Taxes(22%*$13167) $2896.74
Net income $10270.26
Less:Dividends($10270.26*0.377400377) $3876
Addition to retained earnings $6394.26

Total assets would be=$70800*1.14=$80712

Total equity would be=$36200+Addition to retained earnings

=$36200+$6394.26=$42594.26

Total assets=Total equity+Total liabilities

Hence external financing needed=$80712-$42594.26-$34600

which is equal to

=$3517.74